# General task optimal planning for heterogeneous teams with precedence and compatibility constraints and its application on power grid inspection using unmanned aerial vehicles

### v1

@@ 3 Power grid inspection
One of the main application of the TSP and its variants is the vehicle routing problem (VRP) and specifically for UAVs. In this context, several generalization of the TSP and VRP have been used to compute the planification for the inspection of power grids[17, 18, 19]. Even though the cited papers show clear success, they lack the flexibility needed to model some of the most crucial aspects of real-life scenarios. Such aspects include the consideration of the physical limitations of fixed-wing aircrafts for the pylon or tower inspection, the possibility of simultaneous planification for power line segments and tower inspection within the same mixed mission as separate tasks or the realistic modelling of UAV autonomy. The last one is critical as they can render planification useless if not correctly done. Current formulations usually use a simplified version of battery behaviour where only the total energy consumption per UAV is computed in the planification. This is not an issue for short missions where the autonomy is enough to securely power through the complete mission. However, it is not valid for cases where mid-route recharging might be needed. Two approaches are possible to solve this. The first one is to solve the problem iteratively, checking whether a solution correctly recharges batteries before depletion and discarding it if not. The second option is more involved but more effective in both potential applications and computational time. The approach used in this work to compute partial costs within the MILP can be extended to non-monotonic functions and used to keep track of the battery level at each route step. Then a constraint on each vertex is added to forbid negative battery levels.
The HMWTPP formulation proposed at section 2 could be used to solve many of these limitations while keeping a comparable level of numerical performance [^4] We warn however that the focus of this paper is not computational performance and although it is briefly discussed, no much detail will be given. No many numerical optimizations are used and many computations are done by solving the raw MILP without any pre-processing external to the SCIP solver.. Let’s address its implementation in power grid inspection.
@@
Each UAV can be represented as a member of the team of workers whose abstract bases are given by their physical landing stations. The autonomy, speeds and physical characteristics and limitations can be easily encoded into the weights and task compatibility of the weighted directed multigraph. We mainly consider two types of aircrafts: multi-rotor such as the DJI Matrice 300 and fixed-wing (including VTOLs) as the DeltaQuad Pro V2 VTOL (see 2).
The tasked vertices will correspond to each of the towers and cable segments selected for inspection. UAVs would need to complete a whole circular orbit at a fixed distance around each tower and follow along the length of each cable segment. For each tower, one vertex is added to include the corresponding inspection task. For cable segments, two vertices are needed for the two possible approaches for each cable and direction.
The edges follow the general rules already discussed in section 2, where the only incompatibility we consider is that VTOL aircrafts can not inspect towers due to their limited turning radius when in airplane mode.
@@
For multi-rotors, the energy consumption model considered is the one proposed at [21] assuming that all the movements between two waypoints are done in a straight line at either navigation or inspection speed. However, in some cases the eudem25[22] elevation dataset is used to check for possible terrain collision and to modify the path to avoid it (see 3(a)). Usually, this is done by climbing to a safe altitude when moving between inspection tasks and then descend to the inspection altitude once the next waypoint is reached. This, of course adds up to the energy and time costs.
For fixed-wing and VTOLs, the same energy model is used while an specific model for such type is being developed by our research group. It is clear that the results obtained after this decision can be non-sensical, however the main points and reasoning remain the same. The trajectories for this type of UAV are computed using the optimal Dubins’ paths [23] between two waypoints. Dubins’ path take into account the limited turning radius to obtain the shortest path between to points with some specific entry and exit directions. These directions in our case are given by the directions of the line segment inspected or by the wind direction in case of take off or landing (see 3(b)). This wind data is retrieved pre-planification using any of the weather data APIs available, in our case, the OpenWeather API[24]. With these types of aircrafts, the mission starts with a take off againts the wind direction followed by a climb to a safe altitude usually higher than the one for multi-rotors. Possible collisions with terrain are also checked and the path modified if needed.
@@
Figure 3:
@@
are also added to check whether the UAVs’ autonomies are enough for the entire mission execution. They also help with numerical performance as the feasible solution space is reduced.
Consider however the case where no possible route exists to complete the inspection due to autonomy issues. In real applications, two options might be considered to circumvent this. The first option is to segment the inspection into sequences where the team inspect a subset of the structures, then gets back to base to recharge and then continues. However, a more efficient but complex solution exists. A set of charging stations can be distributed through the power grid to allow for autonomous recharge. The decision of when to pause the inspection and recharge can be taken either manually or by the solver if we introduce some new constraints to keep track of the battery level at each step of the route in a similar fashion to the MFE constraints $\mathcal{C}_{\text{MFE}}$ Equation 19, Equation 20, Equation 21 and Equation 22555Notice however that these are not directly valid as battery levels do not have monotonic behaviour if charging stations are used. By adding new non-mandatory tasked vertices to $V$ we can represent the charging stations and process, and by defining the energy costs associated to them as negative, we allow the model to manage autonomous recharging. The amount of charge and time spent at each station can be also be a free parameter in the MILP similar to the wait time defined at subsection 2.4 or be fixed beforehand.
Once everything is set up and the MILP is solved, the solution is then parsed into a sequences of physical waypoints based on the original paths used to compute the time and energy weights to control the speed, location and the camera gimbal angle (if available) of each UAV. This data is then sent to the ground control station (GCS) and then sent to the corresponding autopilot. In our case, the GCS was developed by [26].
@@
Figure 4: The ATLAS power grid used for testing. The red dot represent the two bases used. The blue dots represent the pylons/towers and the black lines the power cables between them.
@@
The main results are mostly similar to those obtained with the TSPLIB instances in subsection 2.3. However some extra nuances appear as we are now using different UAVs in multi-UAV missions and wind is present. In this application, the use of the iterative solving method using DFJ SECs is not that effective. The convergence for $n_{\text{U}}\geq 3$, i.e., the number of UAVs, is too slow and renders its usage ineffective. It is clear that thew use of MTZ SECs is the superior choice, not only yielding solution faster but also providing the necessary framework to introduce order and precedence constraints.
Another interesting behaviour is the one obtained when solving the same instance with different UAV teams and wind direction and speed. In general, the less symmetric the problem, the faster its solution is found. This can be seen by solving U4T13 several times, each with a more diverse UAV team and fixed wind speed at $[2,2,0]$ m/s. All the UAVs start at B00. The results can be seen in Table 3.
@@
However, due to the frequent appearance of long branches in the topology of power grids, clustering can be used to improve performance while not loosing much information nor detail about the problem. In this context, clustering refers to the association of several physical structures into a single inspection task. As the optimal routes usually follow an entire branch or a big fraction of it from its initial point until the end, a natural clustering rule would be to cluster all the towers and line segments into a single tasked vertex using some kind of heuristics. This would reduce the number of total vertices in the problem thus decreasing the computation time and allowing for the easy access to the solution of large-scale problems. Promising results have already being achieved and our own research group has develop a simple yet effective clustering algorithm [27].
As a last test, a very simple instance with $4$ towers and $5$ UAVs is solved. All UAVs are the same except for a slower one. In this case, the model correctly identified and deactivated the slower one, assigning one tower inspection to each of the other UAVs.
@@ 4 Conclusion
In this paper we address the unification of many of the TSP-based problems used to model the planification of task execution into a single problem. We explicitly show that our proposed model can be used to compute the optimal route that visits all the points within a graph using several visitors, to compute the optimal plan of autonomous assembly line and to compute the optimal plan for autonomous power grid inspection. It is not difficult to see that it can be also applied in other fields with little to no additional modification. The cover range is wide as it can be used to generalize the genome reconstruction method used in [28] or to compute the Fourier coefficients of the electron density function of a crystal [29] while also providing the ability to compute the optimal route for package delivery with time windows and precedence constraints [30] (and with capacity limit if correctly modified). The proposed brings an effective way to introduce TSP-based models into new problems its flexibility allows for easy extensions and modifications.
This flexibility however comes at a cost: the increase in numerical complexity and poor performance. As the HMWTPP needs to include many different variables to keep track of several per step local parameters, it is a more complex problem than many other TSP-based models. This can be seen in subsection 2.3 where many classical TSPLIB instances have been solved. The time needed to solve these problems are either up to par or several order of magnitude slower than with the use of their TSP and ATSP counterparts and evermore so if heuristic algorithms are used [31, 32, 33]. This means that optimization is needed and it is proposed as future work. Such future work should analyze the use of heuristics and computation of new inequalities to speed up the HMWTPP solving process. It should also explore its extension to include inequalities for non monotonic partial cost functions.

### v2

@@ 3 Power grid inspection
One of the main applications of the TSP and its variants is in the vehicle routing problem (VRP) and specifically for UAVs. In this context, several generalization of the TSP and VRP have been used to compute the planning for the inspection of power grids[20, 21, 22]. Even though the cited papers show clear success, they lack the flexibility needed to model some of the most crucial aspects of real-life scenarios. Such aspects include the consideration of the physical limitations of fixed-wing aircrafts for the pylon or tower inspection, the possibility of simultaneous planning for power line segments and tower inspection as separate tasks within the same mixed mission or the realistic modeling of UAV autonomy. The last one is critical as they can render planning useless if not correctly done. Current formulations usually use a simplified version of battery behavior where only the total energy consumption per UAV is computed in the planning. This is not an issue for short missions where autonomy is enough to securely power through the complete flight. However, it is not valid for cases where mid-route recharging might be needed. Two approaches are possible to solve this. The first one is to add recharging stations vertices, then solve the problem iteratively, checking whether a solution correctly recharges batteries before depletion and discarding it if not. The second option is more involved but more effective in both potential applications and computationally. The approach used in this work to compute partial costs within the MILP can be extended to non-monotonic functions and used to keep track of the battery level at each route step. Then a constraint on each vertex is added to forbid negative battery levels.
The HMWTPP formulation proposed at section 2 could be used to solve many of these limitations while keeping a comparable level of numerical performance [^4] As already said, computational performance is not the focus of this paper and although it is briefly discussed, not much detail is given. Let’s address its implementation in power grid inspection.
@@
Each UAV can be represented as a member of the team of workers whose abstract bases are given by their physical landing stations. The autonomy, speeds and physical characteristics and limitations can be easily encoded into the weights and task compatibility of the WDM graph. We mainly consider two types of aircraft: multi-rotor such as the DJI Matrice 300 and fixed-wing (including VTOLs) such as the DeltaQuad Pro V2 VTOL (see 3).
The tasked vertices will correspond to each of the towers and cable segments selected for inspection. UAVs would need to complete a whole circular orbit at a fixed distance around each tower and follow along the length of each cable segment. For each tower, one vertex is added to include the corresponding inspection task. For cable segments, two vertices are needed for the two possible inspection directions.
The edges follow the general rules already discussed in section 2, where the only incompatibility we consider is that VTOL aircrafts cannot inspect towers due to their limited turning radius when in fixed-wing mode.
@@
For multi-rotors, the energy consumption model considered is the one proposed at [24] assuming that all the movements between two waypoints are done in a straight line at either navigation or inspection speed. However, in some cases the eudem25[25] elevation dataset is used to check for possible terrain collision and to modify the path to avoid it (see 4(a)). Usually, this is done by climbing to a safe altitude when moving between inspection tasks and then descending to the inspection altitude once the next waypoint is reached. This, of course adds up to the energy and time costs.
For fixed-wing and VTOLs, the same energy model is used while a specific model for such type is being developed by our research group. It is clear that the results obtained after this decision can be nonsensical, however the main points and reasoning remain the same. The trajectories for this type of UAV are computed using the optimal Dubins’ paths [26] between two waypoints. Dubins’ paths take into account the limited turning radius to obtain the shortest path between two points with some specific entry and exit directions. These directions in our case are given by the directions of the line segment inspected or by the wind direction in case of takeoff or landing (see 4(b)). This wind data is retrieved pre-planning using any of the weather data APIs available, in our case, the OpenWeather API[27]. With these types of aircraft, the mission starts with a takeoff againt the wind direction followed by a climb to a safe altitude usually higher than the one for multi-rotors. Possible collisions with terrain are also checked and the path modified if needed.
@@
Figure 4: Typical routes for a DJI M300 (to the left) and a DeltaQuad Pro V2 (to the right)
@@
are also added to check whether the UAVs’ autonomies are enough for the entire mission execution. They also help with numerical performance as the feasible solution space is considerably reduced.
Consider however the case where no possible route exists to complete the inspection due to autonomy issues. In real applications, two options might be considered to circumvent this. The first option is to segment the inspection into sequences where the team inspects a subset of the structures, then get back to the base, recharge and then continue. However, a more efficient but complex solution exists. A set of charging stations can be distributed throughout the power grid area to allow autonomous recharging. The decision of when to pause the inspection and recharge can be taken either manually or by the solver if we introduce some new constraints to keep track of the battery level at each step of the route in a similar fashion to the MFE constraints $\mathcal{C}_{\text{MFE}}$ Equation 19, Equation 20, Equation 21 and Equation 22555Notice however that these are not directly valid as battery levels do not have monotonic behavior if charging stations are used. By adding new non-mandatory tasked vertices to $V$ we can represent the charging stations and process, and by defining the energy costs associated to them as negative, we allow the model to manage autonomous recharging. The amount of charge and time spent at each station can also be a free parameter in the MILP similar to the wait time defined at subsection 2.4 or be fixed beforehand. Previous works already do this successfully although in a limited manner [11].
Once everything is set up and the MILP is solved, the solution is then parsed into sequences of physical waypoints based on the original paths used to compute the time and energy weights to control the speed, location and the camera gimbal angle (if available) of each UAV. This data is then sent to the ground control station (GCS) and then sent to the corresponding autopilot. In our case, the GCS was developed by [29] (see Figure 5 for a reference screenshot).
Figure 5: The ground control station displaying the optimal planning for three DJI M300 UAVs inspecting 10 cable segments of a sample power grid near El Rocío, Spain obtained using the HMWTPP model.
@@
Figure 6: The ATLAS power grid used for testing. The red dot represents the two bases used. The blue dots represent the pylons/towers and the black lines the power cables between them.
@@
The main results are mostly similar to those obtained with the TSPLIB instances in subsection 2.3. However, some extra nuances appear as we are now using different UAVs in multi-UAV missions and wind is present. In this application, the use of the iterative solving method using DFJ SECs is not that effective. The convergence for $n_{\text{U}}\geq 3$, i.e., the number of UAVs, is too slow and renders its usage not that effective. It is clear that the use of MTZ SECs is the superior choice in many cases, not only yielding a solution faster but also providing the necessary framework to introduce order and precedence constraints.
Another interesting behavior is the one obtained when solving the same instance with different UAV teams and wind direction and speed. In general, the less symmetric the problem, the faster its solution is found. This can be seen by solving U4T13 several times, each with a more diverse UAV team and fixed wind speed at $[2,2,0]$ m/s. All the UAVs start at B00. The results can be seen in Table 3.
@@
However, due to the frequent appearance of long branches in the topology of power grids, clustering can be used to improve performance while not losing much information nor detail about the problem. In this context, clustering refers to the association of several physical structures into a single inspection task. As the optimal routes usually follow an entire branch or a big fraction of it from its initial point until the end, a natural clustering rule would be to cluster all the towers and line segments into a single tasked vertex using some kind of heuristics. This would reduce the number of total vertices in the problem thus decreasing the computation time and allowing for easy access to the solution of large-scale problems. Promising results have already been achieved and our own research group has develop a simple yet effective clustering algorithm [30].
As a last test, a very simple instance, TOOMANY, with $4$ towers and $5$ UAVs is solved. All UAVs are the same model except for a slower one. In this case, the solver correctly identified and deactivated the slower one, assigning one tower inspection to each of the other UAVs.
This task planner is now part of the autonomous Multi-UAV system developed under the context of the spanish RESISTO Project [31] in conjunction with the aforementioned GCS.
@@ 4 Conclusion
In this paper we address the unification of many of the TSP-based problems used to model the planning of task execution into a single mathematical framework. We explicitly show that our proposed model can be used to compute the optimal route that visits all the points within a graph using several visitors, to compute the optimal plan of autonomous assembly line and to compute the optimal plan for autonomous power grid inspection. It is not difficult to see that it can be also applied in other fields with little to no additional modification. The cover range is wide as it can be used to generalize the genome reconstruction method used in [32] or to compute the Fourier coefficients of the electron density function of a crystal [33] while also providing the ability to compute the optimal route for package delivery with time windows and precedence constraints [34] (and with capacity limit if correctly modified). The proposed brings an effective way to introduce TSP-based models into new problems its flexibility allows for easy extensions and modifications.
This flexibility however, comes at a cost: the increase in numerical complexity and poor performance. As the HMWTPP needs to include many different variables to keep track of several per step local parameters, it is a more complex problem than many other TSP-based models. This can be seen in subsection 2.3 where many classical TSPLIB instances have been solved. The time needed to solve these problems is either up to par or several order of magnitude slower than with the use of their TSP and ATSP counterparts and evermore so if heuristic algorithms are used [35, 36, 37]. This means that optimization is needed and it is proposed as future work. Such future work should analyze the use of heuristics and computation of new inequalities to speed up the HMWTPP solving process. It should also explore its extension to include inequalities for non-monotonic partial cost functions. Another possible extensions for future works are the extension to a mixed Zermelo-TSP [38] approach to allow the model to compute more optimal routes dynamically in very windy environments and the addition of autonomous obstacle avoidance in a UAV Team scenario.
5 Acknowledgements
This work was supported by the R+D+i project TED2021-131716B-C22, funded by MCIN/AEI/10.13039/501100011033 and by the European Union NextGenerationEU/PRTR. This work was also supported by Spanish project RESISTO (2021/C005/00144188) funded by FEDER (Fondo Europeo de Desarrollo Regional) from Ministerio de Asuntos Económicos y Transformación Digital.